OpenStudy (anonymous):

How many pairs of integers (x,y) are there such that 2x2−2xy+y2=225?

OpenStudy (anonymous):

2x2−2xy+y2=225 x^2 + (x-y)^2 =225

OpenStudy (anonymous):

15^2 + 0^2 = 225 =>x = 15 AND x-y=0 Thus, x=15 and y=15

OpenStudy (anonymous):

12^2 + 9^2 =225

OpenStudy (anonymous):

=> x=12 AND x-y=9 Thus, x=12 and y=3 Now, 9^2 + 12^2 =225 =>x=9 and x-y=12 So, x=9 and y=-3

OpenStudy (anonymous):

So, looks likes (15,15) , (12,3) and (9,-3) are the only solutions

OpenStudy (anonymous):

no i think there will be more number of solutions are available for this equation

OpenStudy (anonymous):

(-15,-15) , (-12 , -21) , (-9,-21)

OpenStudy (anonymous):

0^2 + 15^2 =225 0^2 + (-15)^2=225 15^2 + 0^2=225 (-15)^2 + 0^2 =225 (12)^2 + (9)^2=225 (-12)^2+(9)^2=225 (12)^2 + (-9)^2=225 (-12)^2 + (-9)^2=225 (9)^2 + (12)^2 =225 (-9)^2+(12)^2=225 (9)^2+(-12)^2=225 (-9)+(-12)^2=225

OpenStudy (sirm3d):

\[\large x^2+(x-y)^2=225\] and \[\large 9^2+12^2=15^2\] so there are the possibilities:\[\large x=\pm 9 \text { and } x-y = \pm 12\] \[\large x=\pm 12 \text { and } x-y = \pm 9\] and the trivial cases \\large [x=0\] and \[\large x= \pm 15\]

OpenStudy (anonymous):

Looks like 12 solutions

OpenStudy (anonymous):

thank u 12 is correct

OpenStudy (anonymous):

yw